All parts of this multi-part (simple, five-line) flexible budget are based upon the same set of data. Thermaltake Company prepared a budget for an anticipated production level of 50,000 units and an anticipated sales level of 50,000 units. That existing budget for the 50,000 unit level showed lines of $500,000 for sales, $250,000 for variable costs, $250,000 for contribution margin, $100,000 for fixed costs, and $150,000 for net income. Actual production and sales levels experienced during the period were 60,000 units. In the flexible budget that reflects adjustments appropriate for the new actual level of 60,000 units,

1. the appropriate dollar amount for sales should be:

2. the appropriate dollar amount for variable costs should be:

3. the appropriate dollar amount for contribution margin should be:

4. the appropriate dollar amount for fixed costs should be:

5. the appropriate dollar amount for net income should be:
Answer #1

a) Amount of sales = 500000/50000*60000 = $600000

b) Amount of variable cost = 250000/50000*60000 = $300000

c) Contribution margin = 600000-300000 = $300000

d) Fixed cost = 100000

e) Net income = 300000-100000 = 200000